OpenStudy (anonymous):

Simplified form of 3a^4 b^-2 c^3 They are connected in the problem. Also what is the simplified form of -(14x)^0 y^-7 z They are connected in the problem I just had to space them also ^ means exponent Please help!

OpenStudy (anonymous):

\[3a ^{4}b ^{-2}c ^{3}=\frac{ 3a ^{4}c ^{3} }{b ^{2} }\]

OpenStudy (anonymous):

\[-\left( 14x \right)^{0}y ^{-7}z=\frac{ -z }{ y ^{7} }\]